OpenStudy (ny,ny):

Given y= f(t) = 2/t, find the derivative of y with respect to t.

satellite73 (satellite73):

the derivative of \(\frac{1}{x}\) is \(-\frac{1}{x^2}\) always

satellite73 (satellite73):

unless you are supposed to do this by hand

OpenStudy (eliesaab):

If \[ y=t^n\] then \[ y'= n t^{n-1} \]

OpenStudy (eliesaab):

Apply the above for n=-1;

OpenStudy (ny,ny):

umm. (2/t)^-1? I think I am supposed to do this using the limit formula though.

satellite73 (satellite73):

that is what i suspected

OpenStudy (jackthegreatest):



OpenStudy (jackthegreatest):

this correct?

satellite73 (satellite73):

\[\frac{\frac{1}{t+h}-\frac{1}{t}}{h}\]

satellite73 (satellite73):

then a bunch of algebra, subtract first up top

OpenStudy (ny,ny):

yes that is the formula. @satellite73 i know the answer to this question so what you said earlier about 1/x makes sense. I just don't know how to do this, and the teacher used the formula to solve this. and ok, 2t-2(t+h) ---------- t(t+h) ?

satellite73 (satellite73):

sure looks good

satellite73 (satellite73):

numerator is just \(-2h\) when you distribute and combine like terms

satellite73 (satellite73):

so you are looking at \[\frac{-2h}{t(t+h)}\] now divide by \(h\)

OpenStudy (ny,ny):

-2h/th(t+h) ?

satellite73 (satellite73):

cancel the h top and bottom

satellite73 (satellite73):

gives \[\frac{-2}{t(t+h)}\] then put \(h=0\) to get the answer you know it has to be

OpenStudy (ny,ny):

ohhh. oh i forgot about substituting. ok. -2/t(t+0) = -2/t^2 thank you so much.

satellite73 (satellite73):

yw